# Qini Curve Estimation Under Clustered Network Interference

### v1

@@ 6.2 How does interference affect the estimation error?
In the first experiment, we evaluated all estimation strategies under different interference structures by varying $\eta$ as described in the previous subsection. In addition, we fixed the number of buyers (i.e., clusters) while varying the number of items $M$ (i.e., units) as the spillover effects due to interference largely is expected to depend on the cluster size; when the cluster size equals one, there is no interference.
We compared the bias and mean squared error (MSE) of each estimation strategy, as shown in Figure 2. Starting with the bias, we observed that the naive estimator is significantly biased for all interference structures. The bias, which increased with number of items, was generally smaller for the fractional IPW and $\beta$-IPW estimators with $\beta=2$. Meanwhile, the standard IPW estimator appeared unbiased in all cases. In all cases, the $\beta$-IPW estimator with $\beta=2$ is observed to have lower bias than the variant with $\beta=1$ and, in some cases, for larger number of items (e.g., more than 10) the variant with $\beta=1$ had a similar or larger absolute bias as the naive estimator.
By examining the Qini curves for a fixed number of items $M=11$, as in Figure 3, we can obtain a qualitative assessment of the bias of each estimation strategy. Most notably, the naive estimator and $\beta$-IPW estimator with $\beta=1$ outputs significantly different Qini curves despite having a similar absolute bias. For the naive estimator, the Qini curves are biased for the higher percentages of treated, whereas the bias of the $\beta$-IPW estimator is more uniformly biased.
Lastly, in terms of MSE, the standard IPW and fractional IPW estimators performed worst, with MSE increasing exponentially with cluster size. In contrast, the $\beta$-IPW estimators performed the best. In particular, for the product function, $\beta=1$ yielded the best MSE. The naive estimator performed the worst for small item counts but its increase in MSE appeared to slow down for larger number of items.
@@
6.3 Which estimation strategy is most efficient?
In the next experiment, we evaluated the efficiency of each estimation strategy by reporting the variance as we varied the number the buyers $N$ (i.e., clusters) or items $M$ (i.e., units). While varying one, the other was kept fixed to $N=20.000$ and $K=11$. We present results only using $\eta_{\text{exp-decay}}$ as we observed no difference when changing this function.
The results indicate that the most efficient estimators, ranked from lowest to highest variance, are: naive, $\beta$-IPW $(\beta=1)$, $\beta$-IPW $(\beta=2)$, fractional IPW, and standard IPW. The variance of the standard IPW and fractional IPW estimators appeared to increase exponentially with the number of buyers $K$, whereas the others scaled sub-exponentially. Full results are provided in Figure 4 in the appendix.
@@ 6.4 How well can we rank policies under clustered network interference?
In the final experiment, we evaluated each estimation strategy’s ability to rank policies based on the estimated area under the Qini curve. To do so, we degraded the optimal treatment prioritization rule by adding progressively larger noise, generating seven policies with decreasing performance. This experiment was repeated 200 times with $N=20.000$ and $K=11$ for each $\eta$, and we report the average Kendall rank correlation coefficient between each strategy’s ranking and the ground truth ranking.
The results, presented in Table 1 in the appendix, show that $\beta$-IPW with $\beta=1$ performed best overall, achieving a correct ranking in all cases (rank correlation coefficient = 1). However, the other estimators also attained similar or only slightly lower rank correlations, except in the case of $\eta_{\text{product}}$ where the naive estimator, standard IPW, and fractional IPW had rank correlations between 0.80 and 0.85.
@@ 7 Discussion
Our findings indicate that, while clustered network interference can cause severe bias in Qini curve estimates, it is possible to get accurate estimates using different estimation strategies that take interference into account. However, the best estimation strategy will depend on several application-specific factors, including cluster size, the number of observations, and prior beliefs about interference and the intended use of the Qini curve.
@@
For small cluster sizes (e.g., fewer than 5), the choice of estimation strategy had a limited impact on estimation error, as IPW, fractional IPW, and $\beta$-IPW all performed comparably in terms of bias and mean squared error. For larger cluster sizes, however, we observed a trade-off between using an unbiased, high-variance estimator and a possibly biased, low-variance estimator. For unbiased estimation, the standard IPW estimator is preferred as it relies only on the weakest conditions regarding the interference, though it requires a large number of observations to be reliable. When data is limited, unbiased estimation might still be feasible if strong domain expertise about the interference structure can justify the use of either the fractional IPW or the $\beta$-IPW estimator. However, if some bias is tolerated, then our results suggest that $\beta$-IPW is a strong choice, as it has the lowest variance among strategies that account for interference. Increasing $\beta$ seems to reduce bias at the cost of introducing more variance; a possibly interesting research direction is to investigate how to data-adaptively tune $\beta$ for this bias-variance trade-off by e.g. fitting a polynomial model based on the $\beta$-additive assumption.
@@
The acceptable level of bias depends on the decision-making context. For model selection –i.e., discriminating between good and bad policies– we observed the $\beta$-IPW estimator performing best. Interestingly, the naive estimator that ignores interference also ranked policies correctly in some cases, suggesting that interference-related bias may have a limited effect on this type of decision-making. However, if the goal is to determine a suitable threshold for a treatment prioritization rule, a well-calibrated Qini curve becomes more critical, making an estimator with low bias preferable.
An additional consideration, which was not explored in this study, is estimating Qini curves under clustered network interference from observational data when the propensity score is unknown. One could replace the known propensity score with the estimated propensity score in all estimators that we discussed, potentially introducing other forms of biases. These may arise from model misspecification, which could be mitigated with doubly-robust estimators (Liu et al., 2019), or from unmeasured confounders, which could be addressed through sensitivity analysis (VanderWeele et al., 2015) to establish bounds under varying degrees of confounding. Future research should investigate these strategies in the context of Qini curve estimation.
@@ 8 Conclusion
We have introduced a framework for estimating Qini curves in experimental study designs with clustered network interference, along with multiple estimation strategies. Our results demonstrate that properly accounting for interference leads to more accurate Qini curve estimation, though the best estimation strategy depends on the specific context. To guide practitioners, we provide practical recommendations based on both theoretical insights and empirical findings, helping them better assess the cost-effectiveness of treatment policies in complex settings with interference.
@@ Acknowledgments
This work was primarily conducted during an internship of RKAK at Booking.com. We appreciate the feedback from discussions by colleagues and would like to thank Mathijs de Jong, Ilir Maçi, Alina Solovjova, and Antonio Castelli for their feedback which has greatly improved the quality of our work.
@@ Appendix A Proofs and derivations
Under assumption 3.1, we can show that $\mathbb{E}[\phi(\pi)]=V(\pi)$ by rewriting the expectation as follows,
$\mathbb{E}[\phi(\pi)]$ $=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}\mid\mathbf{W}=\boldsymbol{\pi}_{i},X_{i}]\right]$
$=\frac{1}{N}N\mathbb{E}\left[\mathbb{E}\left[Y_{i}\mid\mathbf{W}=\boldsymbol{\pi}_{i},X_{i}\right]\right]$
@@
A.2 Proof of unbiasedness and variance for IPW estimator
@@
The unbiasedness of $\widehat{\phi}^{\text{\;IPW}}(\pi)$ follows from the same arguments as deriving the inverse probability weighting estimator in settings with no interference [Hernan and Robins, 2020]. We can show that
$\mathbb{E}\left[\widehat{\phi}^{\text{\;IPW}}(\pi)\right]$ $=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})}{\Pr(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\mid X_{i})}Y_{i}\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\frac{\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})}{\Pr(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\mid X_{i})}Y_{i}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[Y_{i}\mid\mathbf{W}_{i}=\boldsymbol{\pi}_{i},X_{i}\right]\right]$
@@
The unbiasedness of $\widehat{\phi}^{\text{\;IPW}}(\pi)$ then follows from $\mathbb{E}\left[\phi(\pi)\right]=\mathrm{Var}(\pi)$ according to lemma 4.1.
@@
$\mathrm{Var}\left(\widehat{\phi}^{\text{\;IPW}}(\pi)\right)=\mathrm{Var}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}\right)=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}\right)~{}.$
@@
$\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}\right)=\mathbb{E}\left[\underbrace{\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right)}_{(a)}\right]+\mathrm{Var}\left(\underbrace{\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]}_{(b)}\right)~{}.$
@@
$(b)$ $=\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}(\boldsymbol{\pi}_{i})\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]$
$=Y_{i}(\boldsymbol{\pi}_{i})\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]$
$=Y_{i}(\boldsymbol{\pi}_{i})~{}.$
where it follows from conditional exchangeability in assumption 3.1 that $\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]=\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\mid X_{i}\right]=1$.
@@
$(a)$ $=\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}(\boldsymbol{\pi}_{i})\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right)$
$=\left[\frac{Y_{i}(\boldsymbol{\pi}_{i})}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\right]^{2}\mathrm{Var}\left(\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})\mid Y_{i}(\boldsymbol{\pi}_{i}),X_{i}\right)$
$=\left[\frac{Y_{i}(\boldsymbol{\pi}_{i})}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\right]^{2}\mathrm{Var}\left(\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})\mid X_{i}\right)$
@@
$(a)$ $=\left[\frac{Y_{i}(\boldsymbol{\pi}_{i})}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\right]^{2}\left[\prod_{j=1}^{M_{i}}\left\{e_{1}(X_{i})e_{0}(X_{i})+e_{\pi_{ij}}(X_{i})^{2}\right\}-\prod_{j=1}^{M_{i}}e_{\pi_{ij}}(X_{i})^{2}\right]$
$=\left[Y_{i}(\boldsymbol{\pi}_{i})\right]^{2}\left[\prod_{j=1}^{M_{i}}\left\{\frac{e_{1}(X_{i})e_{0}(X_{i})}{e_{\pi_{ij}}\left(X_{i}\right)^{2}}+1\right\}-1\right]$
@@
$\mathrm{Var}\left(\widehat{\phi}^{\text{\;IPW}}(\pi)\right)=\frac{1}{N^{2}}\sum_{i=1}^{M_{i}}\left\{\mathbb{E}\left[\left[Y_{i}(\boldsymbol{\pi}_{i})\right]^{2}\left[\prod_{j=1}^{M_{i}}\left(\frac{e_{1}(X_{i})e_{0}(X_{i})}{e_{\pi_{ij}}\left(X_{i}\right)^{2}}+1\right)-1\right]\right]+\mathrm{Var}\left(Y_{i}(\boldsymbol{\pi}_{i})\right)\right\}$

### v2

@@ 5.2 How does interference affect the estimation error?
In the first experiment, we evaluated all estimation strategies under different interference structures by varying $\eta$ as described in the previous subsection. In addition, we fixed the number of buyers (i.e., clusters) while varying the number of items $M$ (i.e., units), since spillover effects due to interference are largely expected to depend on cluster size; when the cluster size is one, there is no interference. We compared the bias and mean squared error (MSE) of each estimation strategy, as shown in Figure 2.
Starting with the bias, we observed that the naive estimator is significantly biased for all interference structures. The bias, which increased with number of items, was generally smaller for the fractional IPW and $\beta$-IPW estimators with $\beta=2$. Meanwhile, the standard IPW estimator appeared unbiased in all cases. In all cases, the $\beta$-IPW estimator with $\beta=2$ is observed to have lower bias than the variant with $\beta=1$ and, in some cases, for larger number of items (e.g., more than 10) the variant with $\beta=1$ had a similar or larger absolute bias as the naive estimator.
In terms of MSE, the standard IPW and fractional IPW estimators performed worst, with MSE increasing exponentially with cluster size. The $\beta$-IPW estimators performed the best; in particular, for the product function, $\beta=1$ yielded the best MSE. The naive estimator performed the worst for small item counts but its increase in MSE appeared to slow down for larger number of items.
Finally, by visually examining the average Qini curve for a fixed number of items $M=11$ with the exponential decay function, as in Figure 3, we can obtain a more qualitative assessment of the bias of each estimation strategy. Starting from the left, we observe that the naive estimator vastly overestimates the Qini curve; the standard IPW is centered around the true curve; fractional IPW also overestimates, but to a lesser degree than the naive; and the $\beta$-IPW estimators underestimate the Qini curves, where $\beta=2$ has less bias than $\beta=1$ as expected. We further examined the average difference in the AUC of the estimated Qini curves and the ground-truth ones across all interference structures, and we observe similar trends for the other interference structure (max and product); the full results are provided in Appendix D.5.
5.3 Which estimation strategy is most efficient?
In the next experiment, we evaluated the efficiency of each estimation strategy by reporting the variance as we varied the number the buyers $N$ (i.e., clusters) or items $M$ (i.e., units). While varying one, the other was kept fixed to $N=20.000$ and $K=11$. We present results only using $\eta_{\text{exp-decay}}$ as we observed no difference when changing this function. The results are shown in Figure 4 in the appendix, and we summarize our main findings below.
The results indicate that the most efficient estimators, ranked from lowest to highest variance, are: naive, $\beta$-IPW $(\beta=1)$, $\beta$-IPW $(\beta=2)$, fractional IPW, and standard IPW. While the naive has the lowest variance, we recall that it also has the largest bias since it fails to take into account the interference. As expected, the variance of all estimators improved with more buyers (i.e., clusters) at a similar rate. Meanwhile, the variance of the standard and fractional IPW estimators increased exponentially with the number of items (i.e., cluster size), whereas the others scaled sub-exponentially.
@@ 5.3 Which estimation strategy is most efficient?
Table 1: Comparison of ability to rank policies by each estimation strategy. We used $N=20.000$ number of buyers (i.e., clusters) and $K=11$ items (i.e., units) per buyer. We report the average Kendall rank correlation with respect to the ground truth ranking over 150 repetitions. Higher is better, where 1 corresponds to a perfect rank correlation.
Max Product Exponential decay
Estimation strategy
Naive (ignore interference) 1.000 0.808 1.000
Standard IPW 0.928 0.845 0.945
Fractional IPW 0.991 0.806 0.995
$\beta$-IPW ($\beta=1$) 1.000 1.000 1.000
$\beta$-IPW ($\beta=2$) 1.000 0.995 1.000
@@ 5.4 How well can we rank policies under clustered network interference?
In the final experiment, we evaluated each estimation strategy’s ability to rank policies based on the estimated area under the Qini curve. To do so, we degraded the baseline treatment policy used for evaluation by adding progressively larger noise to its treatment prioritization rule, generating seven policies with decreasing performance. This experiment was repeated 150 times with $N=20.000$ and $K=11$ for each $\eta$, and we report the average Kendall rank correlation coefficient between each strategy’s ranking and the ground truth.
The results, shown in Table 1, indicate that $\beta$-IPW with $\beta=1$ performed best overall, achieving perfect rankings in all cases (rank correlation = 1). The other estimators achieved similar or only slightly lower correlations, except for $\eta_{\text{product}}$, where the naive estimator, standard IPW, and fractional IPW had correlations between 0.80 and 0.85.
@@ 6 Discussion
Our findings indicate that, while clustered network interference can cause severe bias in Qini curve estimates when the interference is ignored, it is possible to get accurate estimates using different estimation strategies that take this interference into account. However, the best estimation strategy will depend on several application-specific factors, including cluster size, the number of observations, and prior beliefs about interference and the intended use of the Qini curve.
@@
For small cluster sizes (e.g., fewer than 5), the choice of estimation strategy had a limited impact on estimation error, as IPW, fractional IPW, and $\beta$-IPW all performed comparably in terms of bias and mean squared error. For larger cluster sizes, however, we observed a trade-off between using an unbiased, high-variance estimator and a possibly biased, low-variance estimator. For unbiased estimation, the standard IPW estimator is preferred as it relies only on the weakest conditions regarding the interference, though it requires a large number of observations to be reliable. When the number of observations is limited, low-variance unbiased estimation might still be feasible if strong domain expertise about the interference structure can justify the use of either the fractional IPW or the $\beta$-IPW estimator. However, if some bias is tolerated, then our results suggest that $\beta$-IPW is a strong choice, as it has the lowest variance among strategies that account for interference.
@@
The acceptable level of bias depends on the decision-making context. For model selection – i.e., discriminating between good and bad policies – we observed the $\beta$-IPW estimator performing best. Interestingly, the naive estimator that ignores interference also ranked policies correctly in some cases, suggesting that bias due to interference may have a limited effect on this type of decision-making. However, if the goal is to determine a suitable threshold for a treatment prioritization rule, a well-calibrated Qini curve becomes more critical, making an estimator with low bias preferable.
Additional considerations when using $\beta$-IPW estimator
Since Assumptions 2 and 3 are untestable, it is difficult to choose between the fractional and $\beta$-IPW estimators without domain expertise. However, for the $\beta$-IPW in particular, an alternative could be to adopt a data-adaptive strategy to find an estimator optimal with respect to the MSE. We can leverage the fact that larger values of $\beta$ make the $\beta$-additivity assumption less restrictive. Recall that as $\beta$ increases, the $\beta$-IPW estimator converges to the standard IPW estimator. This relationship can be exploited to estimate the bias for a given $\beta$ by comparing the $\beta$-IPW estimate to that of the unbiased standard IPW estimator. Combining this bias estimate with an estimate of the sampling variance of the $\beta$-IPW estimator obtained via bootstrapping allows us to compute an MSE estimate for each $\beta$, from which we can select the value that minimizes the estimated MSE. A potential limitation of this approach, however, is that the MSE estimate itself may have high variance, owing to the uncertainty propagated from the standard IPW estimator used as a reference for estimating the bias. Despite this practical challenge, we view this approach as a promising direction for future work toward developing more data-adaptive methods for choosing among the estimators proposed here.
@@ 7 Conclusion
To summarize, we have introduced a framework for estimating Qini curves in experimental study designs with clustered network interference, along with multiple estimation strategies. Our results demonstrate that properly accounting for interference leads to more accurate Qini curve estimation, though the best estimation strategy depends on the specific context. While these methods are not a universal solution for all types of interference and should be applied with care, especially in high-stakes settings, we provide practical recommendations based on both theoretical insights and empirical evidence. These guidelines are intended to help practitioners more effectively assess the cost-effectiveness of treatment policies in complex settings where interference is present.
@@ Acknowledgments
This work was primarily conducted during an internship of RK at Booking.com. We appreciate the feedback from discussions by colleagues and internal reviewers which greatly improved earlier versions of this manuscript.
@@ A.1 Proof of Lemma 1
Under Assumption 1, we can show that $\mathbb{E}[\phi(\pi)]=V(\pi)$ by rewriting the expectation as follows,
$\phi(\pi)$ $=\mathbb{E}\left[\mathbb{E}[\tilde{Y}\mid\mathbf{W}=\boldsymbol{\pi},X]\right]$
@@
A.2 Proof of Theorem 1
@@
The unbiasedness of the standard IPW estimator $\widehat{\phi}^{\text{\;IPW}}(\pi)$ follows from the same arguments as deriving the inverse probability weighting estimator in settings with no interference (Hernan & Robins, 2020). We can show that
$\mathbb{E}\left[\widehat{\phi}^{\text{\;IPW}}(\pi)\right]$ $=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})}{\Pr(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\mid X_{i})}\tilde{Y}_{i}\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\frac{\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})}{\Pr(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\mid X_{i})}\tilde{Y}_{i}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\tilde{Y}_{i}\mid\mathbf{W}_{i}=\boldsymbol{\pi}_{i},X_{i}\right]\right]$
@@
The unbiasedness of $\widehat{\phi}^{\text{\;IPW}}(\pi)$ then follows from $\mathbb{E}\left[\phi(\pi)\right]=V(\pi)$ under Assumption 1.
@@
$\mathrm{Var}\left(\widehat{\phi}^{\text{\;IPW}}(\pi)\right)=\mathrm{Var}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}\right)=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}\right)~.$
@@
$\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}\right)=\mathbb{E}\left[\underbrace{\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right)}_{(a)}\right]+\mathrm{Var}\left(\underbrace{\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]}_{(b)}\right)~.$
@@
$(b)$ $=\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]$
$=\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]$
$=\tilde{Y}_{i}(\boldsymbol{\pi}_{i})~.$
where it follows from conditional exchangeability in Assumption 1 that $\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right]=\mathbb{E}\left[\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\mid X_{i}\right]=1$.
@@
$(a)$ $=\mathrm{Var}\left(\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right)$
$=\left[\frac{\tilde{Y}_{i}(\boldsymbol{\pi}_{i})}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\right]^{2}\mathrm{Var}\left(\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})\mid\tilde{Y}_{i}(\boldsymbol{\pi}_{i}),X_{i}\right)$
$=\left[\frac{\tilde{Y}_{i}(\boldsymbol{\pi}_{i})}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\right]^{2}\mathrm{Var}\left(\mathbf{1}(\mathbf{W}_{i}=\boldsymbol{\pi}_{i})\mid X_{i}\right)$
@@
$(a)$ $=\left[\frac{\tilde{Y}_{i}(\boldsymbol{\pi}_{i})}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\right]^{2}\left[\prod_{j=1}^{M_{i}}\left\{e_{1}(X_{i})e_{0}(X_{i})+e_{\pi_{ij}}(X_{i})^{2}\right\}-\prod_{j=1}^{M_{i}}e_{\pi_{ij}}(X_{i})^{2}\right]$
$=\left[\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\right]^{2}\left[\prod_{j=1}^{M_{i}}\left\{\frac{e_{1}(X_{i})e_{0}(X_{i})}{e_{\pi_{ij}}\left(X_{i}\right)^{2}}+1\right\}-1\right]$
@@
$\mathrm{Var}\left(\widehat{\phi}^{\text{\;IPW}}(\pi)\right)=\frac{1}{N^{2}}\sum_{i=1}^{M_{i}}\left\{\mathbb{E}\left[\left[\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\right]^{2}\left[\prod_{j=1}^{M_{i}}\left(\frac{e_{1}(X_{i})e_{0}(X_{i})}{e_{\pi_{ij}}\left(X_{i}\right)^{2}}+1\right)-1\right]\right]+\mathrm{Var}\left(\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\right)\right\}$